One number is 6 more than another. If the sum of the smaller number and 3 times the larger number is 34, find the two numbers.
Answer by funmath(2933)

One number is 6 more than another. If the sum of the smaller number and 3 times the larger number is 34, find the two numbers.
Let the smaller number be: x
Then the larger number is 6 more than x: x+6
sum means +
3times means 3*
is means =
The problem to solve is:
smaller number + 3* larger number=34
(x)+3(x+6)=34
x+3x+18=34
4x+18=34
4x+18-18=34-18
4x=16
4x/4=16/4
x=4
The smaller number is: x=4
The larger number is: x+6=4+6=10
:
Check to see if this is reasonable, does the sum of the smaller number and three times the larger number=34?
(4)+3*(10)=34
4+30=34
34=34 It checks out.
Happy Calculating!!!